# Movable Antenna Empowered Covert Dual-Functional Radar-Communication

### v1

@@ IV-A Problem Reformulation
IV-B Updating Auxiliary Variable
@@
IV-C Updating Transmit Beamforming
@@
IV-D Updating Transmit Antenna Placement
@@
where $\bm{t}^{l}$ is the obtained APV in the $l$-th iteration, $\tilde{\Gamma}_{w,i}(\bm{t})=\bm{u}_{w,i}^{H}(|\alpha_{w}|^{2}\bm{A}_{w}(\bm{r},\bm{t})\bm{R}_{X}^{i}\bm{A}_{w}(\bm{r},\bm{t})^{H}-\sum_{c\neq w}^{W}\Gamma_{w}|\alpha_{c}|^{2}\bm{A}_{c}(\bm{r},\bm{t})\bm{R}_{X}^{i}\bm{A}_{c}(\bm{r},\bm{t})^{H}-\sigma_{r}^{2}\Gamma_{w}\bm{I}_{N})\bm{u}_{w,i}$, and $\mathcal{G}_{w}(\bm{t})=\bm{a}_{t}^{H}(\varphi_{w},\bm{t})\left(|\beta_{w}|^{2}\bm{R}_{X}^{1}-\kappa_{w}|\beta_{w}|^{2}\bm{R}_{X}^{0}\right)\bm{a}_{t}(\varphi_{w},\bm{t})+(1-\kappa_{w})\sigma_{0}^{2}.$ Note that the positive real numbers $\tilde{\delta}_{w,i}$ and $\bar{\delta}_{w}$ are selected to satisfy $\tilde{\delta}_{w,i}\bm{I}_{N}\succeq\nabla^{2}\tilde{\Gamma}_{w,i}(\bm{t})$ and $\bar{\delta}_{w}\bm{I}_{N}\succeq\nabla^{2}\mathcal{G}_{w}(\bm{t})$, with $\nabla^{2}\tilde{\Gamma}_{w,i}(\bm{t})\in\mathbb{C}^{N\times N}$ and $\nabla^{2}\mathcal{G}_{w}(\bm{t})\in\mathbb{C}^{N\times N}$ being the Hessian matrices, respectively. Please refer to Appendix D for the construction of $\tilde{\delta}_{w,i}$ and $\bar{\delta}_{w}$. Thus, combing (31a) and (31b), the problem in (29b) can be approximated as follows
@@
IV-E Updating Receive Filter and Antenna Placement
@@
which is convex, and can solved by using the CVX tool. Then, $\varrho$ can be updated by
@@
As for $\{\bm{u}_{w,i}\}$, it can be readily observed that the corresponding subproblem for each filter is a generalized Rayleigh quotient maximization problem [1]. The optimal solution $\bm{u}_{w,i}^{\star}$ can be directly given by the eigenvector associated with the largest eigenvalue of the matrix $\left(\sum_{c\neq w}^{W}|\alpha_{c}|^{2}\bm{A}_{c}(\bm{r},\bm{t})\bm{R}_{X}^{i}\bm{A}_{c}(\bm{r},\bm{t})^{H}+\sigma_{r}^{2}\bm{I}_{N}\right)^{-1}|\alpha_{w}|^{2}\\ \bm{A}_{w}(\bm{r},\bm{t})\bm{R}_{X}^{i}\bm{A}_{w}(\bm{r},\bm{t})^{H}$. The overall BCD-based optimization algorithm for non-colluding Willies is summarized in Algorithm 2. The convergence behaviour and complexity analysis are presented in Section V.
@@
V Modification to Colluding Scenarios
@@
V-A Covertness Constraint Reformulation
@@
which is still non-convex w.r.t. $\bm{R}_{0}$, $\bm{W}$, and $\bm{t}$. Fortunately, it is observed that the constraint in (V-A) is formulated as the difference between channel capacities under different hypotheses. Therefore, we rewrite $\mathcal{C}_{1}(\bm{R}_{0},\bm{t})$ and $\mathcal{C}_{2}(\bm{W},\bm{R}_{0},\bm{t})$ into equivalent but more favourable forms, respectively. In terms of $\mathcal{C}_{1}(\bm{R}_{0},\bm{t})$, we adopt the MMSE-based transformation method [43]. Specifically, by introducing an auxiliary matrix $\bm{U}_{1}\in\mathbb{C}^{W\times N}$, the mean square error matrix is given by
@@
Next, for $\mathcal{C}_{2}(\bm{W},\bm{R}_{0},\bm{t})$, we introduce the following lemma, the proof of which can be found in [45].
@@
It is observed that the optimization problem in (47) differs from that in non-colluding scenarios only in the constraint in (46). Therefore, we exhibit modifications on the BCD algorithm. The iteration order of optimization variables in a single loop in Algorithm 2 is adjusted to: $\bm{\rho}\rightarrow\bm{\upsilon}\rightarrow\bm{U}_{1}\rightarrow\bm{P}_{1}\rightarrow\bm{P}_{2}\rightarrow\bm{W}~\text{and}~\bm{R}_{0}\rightarrow\bm{t}\rightarrow\bm{r}\rightarrow\{\bm{u}_{w,i}\}$. Note that the constraint in (46) couples beamforming matrices ($\bm{W}$ and $\bm{R}_{0}$) with the transmit APV $\bm{t}$. Consequently, we focus on the specific modifications for optimizing these variables, while omitting the remaining steps that are identical to those in Section IV for brevity.
V-B Updating Transmit Beamforming
@@
V-C Updating Transmit Antenna Placement
@@
V-D Convergence and Complexity Analysis
@@ VI Simulation Results
In our simulation, we assume that the BS is located at (0, 0) m. The users are randomly distributed in a circle centered at (40, 0) m with a radius of 5 m. The numbers of transmitting and receiving paths are identical, i.e., $L_{k}=L=12,\forall k$. The PRM is given by $\bm{\Sigma}_{k}=\text{diag}\{\sigma_{k,1},\dots,\sigma_{k,L}\}$ with $\sigma_{k,l}\sim\mathcal{CN}\left(0,\frac{c_{k}^{2}}{L}\right)$. Note that $c_{k}^{2}=C_{0}d_{k}^{-a}$ denotes the large-scale path loss, where $C_{0}=-30~\text{dB}$, and the path-loss exponents $\alpha$ for communcation and sensing links are set to be is 3.2 and 2.6, respectively. Unless otherwise specified, the number of sensing targets are set as $W=2$, and the elevation angles are set to be $20\degree$ and $105\degree$, respectively. Other parameters: $K=3,N=6,M=32,\Gamma_{w}=\Gamma=10~\text{dB},\epsilon_{w}=\epsilon_{F}=\epsilon=0.05,\forall w,\lambda=0.1~\text{m},d=\frac{\lambda}{2}$, and $D=14\lambda.$
Figure 2: Convergence behaviour.
@@
We first present the convergence behavior of Algorithm 2 in Fig. 2. It can be observed that the covert sum rate under both colluding and non-colluding Willie scenarios monotonically increases and eventually stabilizes as the iterations progress. For instance, with $N=6$ under colluding Willies, the achievable rate improves from 3.6 bps/Hz to 13.1 bps/Hz, highlighting the effectiveness of the proposed design. In most cases, fewer than 10 iterations are sufficient. Note that the number of antennas $N$ has a minor impact on the convergence rate. This is because many subproblems can be optimally or near-optimally solved (e.g., $\bm{\rho},\bm{\upsilon},\bm{W},\bm{R}_{0},\varrho,\bm{P}_{1},\bm{P}_{2},\bm{U}_{1}$, and $\{\bm{u}_{w,i}\}$), which enables the proposed algorithm to quickly converge to a local optimum.
Fig. 3 illustrates the covert sum rate versus transmission power $P_{t}$. It can be observed that the proposed scheme significantly outperforms both the GAS and FPA benchmarks across the entire range of $P_{t}$. This performance gain is attributed to the active movement of antenna elements, which not only enhances the desired channel conditions for both radar and communication, but also provides additional spatial DoFs to effectively mitigate interference. Furthermore, the results reveal a performance gap between different detection models. Specifically, the sum rate achieved in the non-colluding Willies scenario is strictly higher than that of the colluding counterpart. This is because for colluding Willies, the joint signal processing significantly enhances the detection capability. Consequently, for a given covertness level $\epsilon$, the resulting covertness constraint becomes more stringent, thereby restricting the feasible region of optimization. Another interesting observation is that as $P_{t}$ increases, the performance gap between the proposed scheme in the non-colluding Willie scenario and the upper bound scheme gradually widens (from 1.0 bps/Hz to 3.4 bps/Hz). This is because constraints in (26) limit the discrepancy in Willies’ received signal strength under different hypotheses, which in turn restricts transmit power allocation, leading to the widening performance gap.
@@
We then plot the covert rate versus antenna number $N$ with $P_{t}=10$ dBW in Fig. 7. The results exhibit that the covert sum rate of all schemes increases significantly as $N$ scales up, as a larger $N$ provides more spatial multiplexing gains to balance sensing, communication, and covertness requirements. Notably, the proposed scheme consistently maintains a substantial advantage over the FPA benchmark, especially in the small-$N$ regime, demonstrating that antenna mobility can effectively compensate for limited spatial resources to ensure high-level covertness.
Fig. 8 simulates the effect of Willie number $W$ on the covert sum rate with $N=8$ and $P_{t}=15$ dBW. As $W$ scales up, the performance degradation in colluding scenarios is more pronounced than in non-colluding cases. This highlights that the joint signal processing gain among colluding Willies becomes increasingly significant with the expansion of the monitoring network, further shrinking the feasible region for covert transmission. Nonetheless, the proposed MA-enhanced scheme maintains a more stable performance gain compared with the FPA scheme, demonstrating the effectiveness of flexible antenna movement in enhancing transmission covertness.
VII Conclusion
In this paper, we have investigated a movable antenna-enhanced covert DFRC system. A covert sum rate maximization problem was formulated by jointly designing transmit beamforming vectors, receiving filter, and transceiver antenna placement. Both colluding and non-colluding Willies were considered. Specifically, for non-colluding Willies, we developed an efficient BCD-based algorithm, incorporating SDR, PGD, SCA, and Dinkelbach transformation methods. For colluding Willies, we first derived the minimum DEP under the optimal likelihood ratio test. We subsequently proposed an MMSE-incorporated optimization framework. We further provided a comprehensive complexity analysis for the proposed algorithm. Simulation results showed that our method can significantly improve the covert sum rate, and achieve a satisfactory trade-off between the communication and radar performance compared with existing benchmark schemes. Overall, our work underscores the great potential of MAs, and provides a promising solution for future covert-aware 6G wireless networks.
@@ Appendix A Proof of Theorem 2
where $\bm{X}=[\bm{x}(1),\dots,\bm{x}(M)]\in\mathbb{C}^{N\times M}$, and $\bm{N}_{F}=[\bm{n}_{F}(1),\dots,\bm{n}_{F}(M)]\in\mathbb{C}^{W\times M}$. Since the transmitted signals $\bm{X}$ are assumed to be Gaussian distributed, the PDFs of $\bm{Y}_{F}$ under different hypothesis can be respectively given by
@@ Appendix C Derivation of the Gradient in (30)
Appendix D Construction of $\tilde{\delta}_{w,i}$ and $\bar{\delta}_{w}$ in (31)
We only need to show the construction of $\tilde{\delta}_{w,i}$ since the construction of $\bar{\delta}_{w}$ is similar to that of $\tilde{\delta}_{w,i}$, and is thus omitted for brevity. The Hessian matrix $\nabla^{2}\tilde{\Gamma}_{w,i}(\bm{t})$ w.r.t. $\bm{t}$ is given by
@@ Appendix D Construction of $\tilde{\delta}_{w,i}$ and $\bar{\delta}_{w}$ in (31)
$\sin(\frac{-2\pi}{\lambda}\rho({t}_{j},\varphi_{w})+\frac{2\pi}{\lambda}\rho({t}_{n},\varphi_{w})+\angle\bm{R}_{w,w}^{i}(j,n))\cos\varphi_{w}.$ (64)

### v2

@@ III-A Problem Reformulation
III-B Updating Auxiliary Variable
@@
III-C Updating Transmit Beamforming
@@
III-D Updating Transmit Antenna Placement
@@
where $\bm{t}^{l}$ is the obtained APV in the $l$-th iteration, $\tilde{\Gamma}_{w,i}(\bm{t})=\bm{u}_{w,i}^{H}(|\alpha_{w}|^{2}\bm{A}_{w}(\bm{r},\bm{t})\bm{R}_{X}^{i}\bm{A}_{w}(\bm{r},\bm{t})^{H}-\sum_{c\neq w}^{W}\Gamma_{w}|\alpha_{c}|^{2}\bm{A}_{c}(\bm{r},\bm{t})\bm{R}_{X}^{i}\bm{A}_{c}(\bm{r},\bm{t})^{H}-\sigma_{r}^{2}\Gamma_{w}\bm{I}_{N})\bm{u}_{w,i}$, and $\mathcal{G}_{w}(\bm{t})=\bm{a}_{t}^{H}(\varphi_{w},\bm{t})\left(|\beta_{w}|^{2}\bm{R}_{X}^{1}-\kappa_{w}|\beta_{w}|^{2}\bm{R}_{X}^{0}\right)\bm{a}_{t}(\varphi_{w},\bm{t})+(1-\kappa_{w})\sigma_{0}^{2}.$ Note that the positive real numbers $\tilde{\delta}_{w,i}$ and $\bar{\delta}_{w}$ are selected to satisfy $\tilde{\delta}_{w,i}\bm{I}_{N}\succeq\nabla^{2}\tilde{\Gamma}_{w,i}(\bm{t})$ and $\bar{\delta}_{w}\bm{I}_{N}\succeq\nabla^{2}\mathcal{G}_{w}(\bm{t})$, with $\nabla^{2}\tilde{\Gamma}_{w,i}(\bm{t})\in\mathbb{C}^{N\times N}$ and $\nabla^{2}\mathcal{G}_{w}(\bm{t})\in\mathbb{C}^{N\times N}$ being the Hessian matrices, respectively. Please refer to Appendix D for the construction of $\nabla\tilde{\Gamma}_{w,i}(\bm{t}^{l})$, $\nabla\mathcal{G}_{w}(\bm{t}^{l})$, $\tilde{\delta}_{w,i}$ and $\bar{\delta}_{w}$. Thus, combing (33a) and (33b), the problem in (31b) can be approximated as follows
@@
III-E Updating Receive Filter and Antenna Placement
@@
which is convex, and can be solved by using the CVX tool. Then, $\varrho$ can be updated by
@@
As for $\{\bm{u}_{w,i}\}$, it can be readily observed that the corresponding subproblem for each filter is a generalized Rayleigh quotient maximization problem [1]. The optimal solution $\bm{u}_{w,i}^{\star}$ can be directly given by the eigenvector associated with the largest eigenvalue of the matrix $\left(\sum_{c\neq w}^{W}|\alpha_{c}|^{2}\bm{A}_{c}(\bm{r},\bm{t})\bm{R}_{X}^{i}\bm{A}_{c}(\bm{r},\bm{t})^{H}+\sigma_{r}^{2}\bm{I}_{N}\right)^{-1}|\alpha_{w}|^{2}\\ \bm{A}_{w}(\bm{r},\bm{t})\bm{R}_{X}^{i}\bm{A}_{w}(\bm{r},\bm{t})^{H}$. The overall BCD-based optimization algorithm for non-colluding Willies is summarized in Algorithm 2. The convergence behaviour and complexity analysis are presented in Section IV.
@@
IV Modification to Colluding Scenarios
@@
IV-A Covertness Constraint Reformulation
@@
which is still non-convex w.r.t. $\bm{R}_{0}$, $\bm{W}$, and $\bm{t}$. Fortunately, it is observed that the constraint in (IV-A) is formulated as the difference between channel capacities under different hypotheses. Therefore, we rewrite $\mathcal{C}_{1}(\bm{R}_{0},\bm{t})$ and $\mathcal{C}_{2}(\bm{W},\bm{R}_{0},\bm{t})$ into equivalent but more favourable forms, respectively. In terms of $\mathcal{C}_{1}(\bm{R}_{0},\bm{t})$, we adopt the MMSE-based transformation method [47]. Specifically, by introducing an auxiliary matrix $\bm{U}_{1}\in\mathbb{C}^{W\times N}$, the mean square error matrix is given by
@@
Next, for $\mathcal{C}_{2}(\bm{W},\bm{R}_{0},\bm{t})$, we introduce the following lemma [49].
@@
It is observed that the optimization problem in (49) differs from that in non-colluding scenarios only in the constraint in (48). Therefore, we exhibit modifications on the BCD algorithm. The iteration order of optimization variables in a single loop in Algorithm 2 is adjusted to: $\bm{\rho}\rightarrow\bm{\upsilon}\rightarrow\bm{U}_{1}\rightarrow\bm{P}_{1}\rightarrow\bm{P}_{2}\rightarrow\bm{W}~\text{and}~\bm{R}_{0}\rightarrow\bm{t}\rightarrow\bm{r}\rightarrow\{\bm{u}_{w,i}\}$. Note that the constraint in (48) couples beamforming matrices ($\bm{W}$ and $\bm{R}_{0}$) with the transmit APV $\bm{t}$. Consequently, we focus on the specific modifications for optimizing these variables, while omitting the remaining steps that are identical to those in Section III for brevity.
IV-B Updating Transmit Beamforming
@@
IV-C Updating Transmit Antenna Placement
@@
IV-D Convergence and Complexity Analysis
@@ V Simulation Results
In our simulation, we assume that the BS is located at (0, 0) m. The users are randomly distributed in a circle centered at (40, 0) m with a radius of 5 m. The numbers of transmitting and receiving paths are identical, i.e., $L_{k}=L=12,\forall k$. The PRM is given by $\bm{\Sigma}_{k}=\text{diag}\{\sigma_{k,1},\dots,\sigma_{k,L}\}$ with $\sigma_{k,l}\sim\mathcal{CN}\left(0,\frac{c_{k}^{2}}{L}\right)$. Note that $c_{k}^{2}=C_{0}d_{k}^{-a}$ denotes the large-scale path loss, where $C_{0}=-30~\text{dB}$, and the path-loss exponents $\alpha$ for communcation and sensing links are set to 3.2 and 2.6, respectively. Unless otherwise specified, the number of sensing targets are set as $W=2$, and the elevation angles are set to be $20\degree$ and $105\degree$, respectively. Other parameters: $K=3,N=6,M=32,\Gamma_{w}=\Gamma=10~\text{dB},\epsilon_{w}=\epsilon_{F}=\epsilon=0.05,\forall w,\lambda=0.1~\text{m},d=\frac{\lambda}{2}$, and $D=14\lambda.$
Figure 2: Convergence behaviour of the BCD algorithm.
@@
We first present the convergence behavior of Algorithm 2 in Fig. 2. It can be observed that the covert sum rate under both colluding and non-colluding Willie scenarios monotonically increases and eventually stabilizes as the iterations progress. For instance, with $N=6$ under colluding Willies, the achievable rate improves from 3.6 bps/Hz to 13.1 bps/Hz, highlighting the effectiveness of the proposed design. In most cases, fewer than 10 iterations are sufficient. Note that the number of antennas $N$ has a minor impact on the convergence rate. This is because many subproblems can be optimally or near-optimally solved (e.g., $\bm{\rho},\bm{\upsilon},\bm{W},\bm{R}_{0},\varrho,\bm{P}_{1},\bm{P}_{2},\bm{U}_{1}$, and $\{\bm{u}_{w,i}\}$), which enables the proposed algorithm to quickly converge to a local optimum. We further define colluding gain as the performance gap between non-colluding and colluding Willie scenarios. It can be observed that the colluding gain diminishes as the number of MAs increases, falling from 5.08 bps/Hz to 2.11 bps/Hz. This is because the additional spatial DoFs provided by more MAs can be effectively leveraged to counteract the enhanced joint detection capability of colluding Willies.
Fig. 3 illustrates the covert sum rate versus transmission power $P_{t}$. It can be observed that the proposed scheme significantly outperforms both the GAS and FPA benchmarks across the entire range of $P_{t}$. This performance gain is attributed to the active movement of antenna elements, which not only enhances the desired channel conditions for both radar and communication, but also provides additional spatial DoFs to effectively mitigate interference. Furthermore, the results reveal a notable colluding gain. Specifically, the sum rate achieved in the non-colluding Willies scenario is strictly higher than that of the colluding counterpart. This is because for colluding Willies, the joint signal processing significantly enhances the detection capability. Consequently, for a given covertness level $\epsilon$, the resulting covertness constraint becomes more stringent, thereby restricting the feasible region of optimization. Another interesting observation is that as $P_{t}$ increases, the performance gap between the proposed scheme in the non-colluding Willie scenario and the upper bound scheme gradually widens (from 1.0 bps/Hz to 3.4 bps/Hz). This is because constraints in (28) limit the discrepancy in Willies’ received signal strength under different hypotheses, which in turn restricts transmit power allocation, leading to the widening performance gap.
@@
We then plot the covert rate versus antenna number $N$ with $P_{t}=10$ dBW in Fig. 7. The results exhibit that the covert sum rate of all schemes increases significantly as $N$ scales up, as a larger $N$ provides more spatial multiplexing gains to balance sensing, communication, and covertness requirements. It is observed that the colluding gain becomes less pronounced as $N$ increases, which is consistent with the observations in Fig. 2. Notably, the proposed scheme consistently maintains a substantial advantage over the FPA benchmark, especially in the small-$N$ regime, demonstrating that antenna mobility can effectively compensate for limited spatial resources to ensure high-level covertness.
Fig. 8 simulates the effect of Willie number $W$ on the covert sum rate with $N=8$ and $P_{t}=15$ dBW. When $W\leq 4$, the colluding gain remains below $2$ bps/Hz. However, this gap nearly doubles as $W$ increases to $5$, which validates the potent joint detection capability of colluding Willies. Nonetheless, the proposed MA-enhanced scheme maintains a more stable performance gain compared with the FPA scheme, demonstrating the effectiveness of flexible antenna movement in enhancing transmission covertness.
VI Conclusion
In this paper, we have investigated a movable antenna-enhanced covert DFRC system. A covert sum rate maximization problem was formulated by jointly designing transmit beamforming vectors, receiving filter, and transceiver antenna placement. Both colluding and non-colluding Willies were considered. Specifically, for non-colluding Willies, we developed an efficient BCD-based algorithm, incorporating SDR, PGD, SCA, and Dinkelbach transformation methods. For colluding Willies, we first derived the minimum DEP under the optimal likelihood ratio test. We subsequently proposed an MMSE-incorporated optimization framework. We further provided a comprehensive complexity analysis for the proposed algorithm. Simulation results showed that our method can significantly improve the covert sum rate, and achieve a satisfactory trade-off between communication and radar performance compared with existing benchmarks. Overall, our work underscores the great potential of MAs, and provides a promising solution for covert-aware 6G wireless networks.
@@ Appendix A Proof of Theorem 2
where $\bm{X}=[\bm{x}(1),\dots,\bm{x}(M)]\in\mathbb{C}^{N\times M}$, and $\bm{N}_{F}=[\bm{n}_{F}(1),\dots,\bm{n}_{F}(M)]\in\mathbb{C}^{W\times M}$. Since the transmitted signals $\bm{X}$ are assumed to be Gaussian distributed, the PDFs of $\bm{Y}_{F}$ under different hypotheses can be respectively given by
@@ Appendix C Derivation of the Gradient in (32)
Appendix D Construction of approximation in (33)
We only need to show the construction of $\nabla\tilde{\Gamma}_{w,i}(\bm{t})$ and $\tilde{\delta}_{w,i}$, since the construction of $\nabla\mathcal{G}_{w}(\bm{t})$ and $\bar{\delta}_{w}$ can be performed in a similar fashion, and is thus omitted for brevity. The Hessian matrix $\nabla^{2}\tilde{\Gamma}_{w,i}(\bm{t})$ w.r.t. $\bm{t}$ is given by
@@ Appendix D Construction of approximation in (33)
$\sin(\frac{-2\pi}{\lambda}\rho({t}_{j},\varphi_{w})+\frac{2\pi}{\lambda}\rho({t}_{n},\varphi_{w})+\angle\bm{R}_{w,w}^{i}(j,n))\cos\varphi_{w}.$